OpenStudy (anonymous):

square root of 18x^5

OpenStudy (anonymous):

\[\sqrt{18x ^{5}}\]

OpenStudy (jhannybean):

\[\large \sqrt{18x^5} = (18x^5)^{1/2}\]

OpenStudy (jhannybean):

now we can separate these two... \[\large 18^{1/2} * (x^5)^{1/2}\]

OpenStudy (jhannybean):

\[\large 18^{1/2} = \sqrt{18} = \sqrt{\color{green}{3*3}*2} = 3\sqrt{2}\]\[\large (x^5)^{1/2}= x^{5* 1/2}= ?\]

OpenStudy (anonymous):

9x^1.5

OpenStudy (jhannybean):

what does 5 * 1/2 =?